# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v3

@@ Abstract.
Code generation with large language models (LLMs) is highly sensitive to token selection during decoding, particularly at decision points where uncertainty strongly affects program correctness. Conventional strategies such as greedy decoding treat all tokens uniformly and fail to capture the uncertainty characteristics unique to code, often resulting in suboptimal outputs. In this work, we conduct an empirical analysis and show that a large fraction of generation errors arises from token misranking at high-uncertainty positions, where the correct token is available but not prioritized.
To address this, we introduce AdaDec, an adaptive decoding framework that employs a lookahead-based, uncertainty-aware pause-and-rerank mechanism. AdaDec automatically learns model-specific uncertainty thresholds and selectively invokes reranking when high uncertainty is detected, leveraging lookahead to refine token choice. Across HumanEval+, MBPP+, and DevEval benchmarks, AdaDec yields substantial improvements, achieving up to 20.9% absolute gains in Pass@1 accuracy compared with greedy decoding, while consistently outperforming prior adaptive decoding approaches such as AdapT. Furthermore, by applying reranking only when necessary, AdaDec reduces computational overhead and latency, enhancing efficiency alongside reliability. These findings underscore the value of uncertainty-guided decoding strategies in advancing the robustness and practicality of LLM-based code generation.
@@ 1. Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024). Models such as CodeLlama (Rozière et al., 2024) and DeepSeek-Coder (Guo et al., 2024)—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation, ultimately reducing developer effort and accelerating the software development process (Jiang et al., 2024; Huynh and Lin, 2025).
At the core of code generation with large language models (LLMs) lies the decoding process, which translates learned latent representations into discrete, executable token sequences (Chen et al., 2021a). Decoding for code generation, however, faces domain-specific challenges. Code exhibits distinctive patterns of token uncertainty: certain tokens (e.g., identifiers or control keywords at the start of a line) are intrinsically ambiguous because their correct choice depends on future logic, while others (e.g., syntactically constrained keywords or repeated identifiers) are relatively deterministic. Crucially, mistakes at a few high-uncertainty points can yield irreversible semantic errors (for example, selecting “for” versus “return” at line start fundamentally alters control flow). Figure 1 provides an example. Therefore, the choice of decoding strategy—how the next token is selected or processed—plays a decisive role in balancing efficiency, diversity, and reliability.
@@
A range of strategies have been proposed: deterministic greedy decoding (Radford et al., 2019) is simple and efficient but prone to premature commitment to suboptimal tokens; sampling-based approaches such as top-$k$ (Holtzman et al., 2020) and nucleus (top-$p$) (Holtzman et al., 2020) sampling enhance diversity but often introduce syntactic or semantic errors that are unacceptable for code; and uncertainty-based methods aim to exploit token-level difficulty signals. For instance, AdapT (Zhu et al., 2023) estimates uncertainty by computing cross-entropy loss between model predictions and ground truth tokens, finding that tokens with higher loss values tend to cluster at line starts. Based on this pre-identified pattern rather than dynamic signals during decoding, AdapT assigns higher temperatures to line-initial tokens, but its reliance on randomness makes gains unstable and limited. By contrast, Uncert-CoT (Zhu et al., 2025) dynamically measures uncertainty at decoding time, employ Shannon entropy (Shannon, 1948) or probability difference (the gap between the top-1 and top-2 token probabilities) to decide whether a line-initial token is highly uncertain, in which case single-line chain-of-thought reasoning is triggered. While this avoids uniform treatment of all line starts, it overlooks uncertain tokens beyond line boundaries and has shown only marginal improvements in practice. Overall, existing methods either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, and current uncertainty-based decoding approaches have not yet explored which uncertainty signals are most appropriate to guide intervention at decoding time.
@@
To address this gap, we conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder (Guo et al., 2024) and the HumanEval dataset (Chen et al., 2021b), we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding-1: The analysis reveals that many errors stem from local ranking mistakes—where the correct token exists in the candidate set but is not ranked first—thereby causing logic drift. To further investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, recording decoding information and the rank of the ground-truth token at each step. This enables us to statistically identify the locations of logic drift points and analyze the relationship between logic drift and entropy. Finding-2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
@@
Base on these findings, We introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding (e.g., greedy decoding, sampling) by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method (Lu et al., 2021), scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete. In contrast to prior approaches that either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, AdaDec systematically explores Shannon entropy as a token-level uncertainty signal and applies dynamic, token-level intervention across both line-initial and in-line tokens only at high-uncertainty steps, addressing the identified gap and enables targeted correction of the most consequential decision points.
@@
We conduct extensive experiments on the HumanEval+, MBPP+ and DevEval benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 20.9% over Greedy decoding and consistently outperforms state-of-the-art adaptive decoding methods such as AdapT. In practice, AdaDec selectively pauses decoding at a low rate (averaging 7.59%), and the targeted lookahead introduces moderate additional latency, effectively balancing quality and efficiency. Ablation studies confirm that the learned entropy threshold accurately identifies uncertain decoding steps, and empirical comparisons show that using the learned threshold outperforms a fixed threshold, particularly for models where the fixed threshold deviates substantially from the learned value. Overall, AdaDec generalizes well across model scales and datasets, significantly improving code generation accuracy without incurring excessive runtime overhead.
@@ 2.1.1. Models
To ensure the generality of our findings, we evaluate AdaDec on a diverse set of open-source code generation models with varying scales and architectures. Specifically, we include:
  • DeepSeek-Coder-1.3B-Instruct (DS-1.3B), DeepSeek-Coder-6.7B-Instruct (DS-6.7B) (Guo et al., 2024): These models are trained from scratch on 2 trillion tokens comprising 87% code and 13% natural language, employing a fill-in-the-blank objective and a 16K context window to support project-level code completion and infilling tasks.
  • StableCode-Instruct-3B (ST-3B) (Pinnaparaju et al., 2023): A 2.7B parameter decoder-only model fine-tuned using Direct Preference Optimization (DPO), demonstrating state-of-the-art performance among similarly sized models on MultiPL-E and MT-Bench coding benchmarks.
  • Qwen3-0.6B (QW-0.6B), Qwen3-1.7B (QW-1.7B), Qwen3-4B (QW-4B), Qwen3-8B (QW-8B) (Team, 2025): A family of multilingual large language models ranging from 0.6B to 8B parameters, featuring dynamic mode switching between reasoning and fast-response modes, and supporting a wide range of tasks including code generation.
These models span from 0.6B to 8B parameters and are accessed via HuggingFace Transformers. They cover a range of architectures and instruction-tuning schemes, ensuring a diverse evaluation across different model families and sizes.
@@ 2.1.2. Implementation
Our implementation is based on PyTorch and HuggingFace Transformers, and decoding is performed in a batch-free, single-example fashion to avoid interference between instances. All models are run using half-precision (i.e., torch.float16) to accelerate inference and reduce memory usage. The maximum number of generated tokens is capped at 512 for all methods to ensure consistent comparison across models.
@@ 2.2. RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset (Chen et al., 2021b). We begin by analyzing the locations of logic drifts—points during decoding where the generated sequence diverges from the correct solution—which often lead to incorrect program behavior. We then investigate how token-level uncertainty, quantified using Shannon entropy (Shannon, 1948), relates to the occurrence of such drifts.
@@ 2.2.1. Method
For each programming problem in the HumanEval dataset, we use DS-6.7B to generate code snippets using a greedy decoding strategy. We then filter out the generations that successfully pass all test cases. Based on an initial human analysis, we observe that in several cases, the model begins with a correct sequence but introduces a critical error—often in the form of incorrect code structures—at a specific step. At these error points, the correct next token is not missing from the candidate set but is ranked below the top-1 position (commonly second or third), resulting in an irreversible deviation from the correct logic. This suggests that some generation failures are caused by local ranking errors at specific decoding steps, ultimately leading to incorrect code behavior. We refer to such steps as logic drift points during decoding and aim to identify them.
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. A mismatch is flagged when node types differ, operators or constants are inconsistent, or discrepancies occur in control structures. The mismatch is then traced back to its corresponding line in the original code and extracted as a candidate drift point. For each mismatched line, we further determine whether the drift occurs at the beginning of the line and record this information for statistical analysis. Finally, each candidate drift point is reviewed by a human annotator to confirm whether it represents an actual drift point and to annotate the precise token where the drift begins. For example, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
@@ 2.3.1. Method
We apply this next-token prediction procedure to a range of open-source LLMs, including three code-specific models, DS-1.3B, DS-6.7B, and ST-3B, as well as four general-purpose models from the Qwen3 family: QW-0.6B, QW-1.7B, QW-4B, and QW-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, and 129,165 for Qwen3 series.
@@ 2.3.2. Results
Model DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.5249 0.5028 0.4936 0.4877
Figure 4 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 4 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1, QW-0.6B has approximately 15% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW-1.7B) exceed a value of 4.5, and the rest fall in the range of 3.5 to 4.
@@ 3.2.1. Data Collection
To enable threshold learning, we cast the problem as a binary classification task. Each sample corresponds to a decoding step and is labeled based on whether the ground truth token is ranked top-1 in the predicted distribution. The input feature for classification is the entropy $H(\bm{p}_{t})$, which reflects the model’s uncertainty at that step.
We generate training data by conducting ground-truth–guided next-token prediction with the target model LM on the BigCodeBench dataset (Zhuo et al., 2024), while carefully ensuring that the construction of our training data does not overlap with the evaluation benchmarks (HumanEval+, MBPP+, and DevEval+). To mitigate the risk of data contamination, we manually reviewed problem statements and function signatures to check for potential duplicates or near-duplicates between BigCodeBench and the evaluation sets. In this process, the ground truth token is provided as input at each step instead of the model’s own prediction. During this process, we record the model’s predicted probability distribution over the vocabulary at each decoding step and calculate the entropy of the predicted distribution. Based on whether the ground truth token is ranked as the top-1 prediction, we assign a binary label: 1 if it is top-1, and 0 otherwise. Based on this process, we collect between approximately 100,000 and 160,000 positive samples (labeled as 1) and between 17,000 and 23,000 negative samples (labeled as 0) per model. For example, in the case of DS-6.7B, we obtain 161,856 positive samples and 19,023 negative samples.

### v4

@@ Abstract.
The success of code synthesis using large language models (LLMs) depends heavily on navigating critical decision points during the decoding process. Standard uniform strategies, such as greedy decoding, often fall short because they fail to distinguish between deterministic steps and those characterized by high logical ambiguity. Our empirical analysis identifies a recurring failure mode: "logic drift" caused by the model’s inability to correctly rank viable candidates during high-uncertainty intervals, even when the ground-truth token is available.
To resolve this, we present AdaDec, a framework that introduces a selective pause-then-rerank mechanism into the decoding pipeline. Unlike static methods, AdaDec utilizes learned, model-specific entropy thresholds to identify when the model is "confused" and dynamically triggers a lookahead-based evaluation to re-score candidate tokens.
Across benchmarks including HumanEval+, MBPP+, and DevEval, AdaDec achieves significant performance breakthroughs, boosting Pass@1 accuracy by up to 20.9% absolute over greedy decoding. The framework not only surpasses traditional Beam Search and specialized methods like AdapT in terms of reliability but also maintains high inference efficiency by intervening only at the most consequential steps. These results suggest that uncertainty-aware adaptive strategies are key to making LLM-driven code generation both robust and practical.
@@ 1. Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024). Models such as CodeLlama (rozière2024codellamaopenfoundation) and DeepSeek-Coder (Guo et al., 2024)—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation, ultimately reducing developer effort and accelerating the software development process (Jiang et al., 2024; Huynh and Lin, 2025).
At the core of code generation with large language models (LLMs) lies the decoding process, which translates learned latent representations into discrete, executable token sequences (Chen et al., 2021). Decoding for code generation, however, faces domain-specific challenges. Code exhibits distinctive patterns of token uncertainty: certain tokens (e.g., identifiers or control keywords at the start of a line) are intrinsically ambiguous because their correct choice depends on future logic, while others (e.g., syntactically constrained keywords or repeated identifiers) are relatively deterministic. Crucially, mistakes at a few high-uncertainty points can yield irreversible semantic errors (for example, selecting “for” versus “return” at line start fundamentally alters control flow). Figure 1 provides an example. Therefore, the choice of decoding strategy—how the next token is selected or processed—plays a decisive role in balancing efficiency, diversity, and reliability.
@@
A range of strategies have been proposed: deterministic greedy decoding (Radford et al., 2019) is simple and efficient but prone to premature commitment to suboptimal tokens; beam search (Freitag and Al-Onaizan, 2017) improves results by maintaining multiple candidates but incurs high computational overhead; sampling-based approaches such as top-$k$ (Holtzman et al., 2020) and nucleus (top-$p$) (Holtzman et al., 2020) sampling enhance diversity but often introduce syntactic or semantic errors that are unacceptable for code; and uncertainty-based methods aim to exploit token-level difficulty signals. For instance, AdapT (Zhu et al., 2023) estimates uncertainty by computing cross-entropy loss between model predictions and ground truth tokens, finding that tokens with higher loss values tend to cluster at line starts. Based on this pre-identified pattern rather than dynamic signals during decoding, AdapT assigns higher temperatures to line-initial tokens, but its reliance on randomness makes gains unstable and limited. By contrast, Uncert-CoT (Zhu et al., 2025) dynamically measures uncertainty at decoding time, employing Shannon entropy (Shannon, 1948) or probability difference (the gap between the top-1 and top-2 token probabilities) to decide whether a line-initial token is highly uncertain, in which case single-line chain-of-thought reasoning is triggered. While this avoids uniform treatment of all line starts, it overlooks uncertain tokens beyond line boundaries and has shown only marginal improvements in practice. Overall, existing methods either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, and current uncertainty-based decoding approaches have not yet explored which uncertainty signals are most appropriate to guide intervention at decoding time.
@@
To address this gap, we conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder (Guo et al., 2024) and the HumanEval dataset (Chen et al., 2021), we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding 1: The analysis reveals that many errors stem from local ranking mistakes—where the correct token exists in the candidate set but is not ranked first—thereby causing logic drift. To further investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, recording decoding information and the rank of the ground-truth token at each step. This enables us to statistically identify the locations of logic drift points and analyze the relationship between logic drift and entropy. Finding 2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
@@
Based on these findings, we introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding (e.g., greedy decoding, sampling) by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method (Lu et al., 2021), scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete. In contrast to prior approaches that either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, AdaDec systematically explores Shannon entropy as a token-level uncertainty signal and applies dynamic, token-level intervention across both line-initial and in-line tokens only at high-uncertainty steps, addressing the identified gap and enabling targeted correction of the most consequential decision points.
@@
We conduct extensive experiments on the HumanEval+, MBPP+ and DevEval benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 20.9% over Greedy decoding and consistently outperforms state-of-the-art adaptive decoding methods such as AdapT. In practice, AdaDec selectively pauses decoding at a low rate (averaging 6.75%), and the targeted lookahead introduces moderate additional latency, effectively balancing quality and efficiency. Ablation studies confirm that the learned entropy threshold accurately identifies uncertain decoding steps, and empirical comparisons show that using the learned threshold outperforms a fixed threshold, particularly for models where the fixed threshold deviates substantially from the learned value. Overall, AdaDec generalizes well across model scales and datasets, significantly improving code generation accuracy without incurring excessive runtime overhead.
@@ 2.1.1. Models
To ensure the generality of our findings, we conduct our study on a diverse set of open-source models ranging from 0.6B to 8B parameters, accessed via Hugging Face. These encompass various architectures and instruction-tuning schemes. Specifically, we include code-specialized models: DeepSeek-Coder-1.3B/6.7B-Instruct (DS-1.3B, DS-6.7B) (Guo et al., 2024), StableCode-Instruct-3B (ST-3B) (Pinnaparaju et al., 2023), and Qwen2.5-Coder-1.5B/7B-Instruct (QW2.5-1.5B, QW2.5-7B) (Hui et al., 2024). We also include general-purpose models featuring dynamic reasoning capabilities: the Qwen3 family, including the 0.6B (QW3-0.6B), 1.7B (QW3-1.7B), 4B (QW3-4B), and 8B (QW3-8B) variants (Team, 2025).
@@ 2.1.2. Implementation
Our implementation is built with PyTorch and the Hugging Face Transformers library. Decoding is performed without batching (i.e., using a batch size of 1) to ensure independent processing of each instance. The maximum generation length is capped at 1,024 tokens for all methods to enable fair comparison across models.
@@ 2.2. RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset (Chen et al., 2021). We begin by analyzing the locations of logic drifts—points during decoding where the generated sequence diverges from the correct solution—which often lead to incorrect program behavior. We then investigate how token-level uncertainty, quantified using Shannon entropy (Shannon, 1948), relates to the occurrence of such drifts.
@@ 2.2.1. Method
For each programming problem in the HumanEval dataset, we use DS-6.7B to generate code snippets using a greedy decoding strategy. We then filter out the generations that successfully pass all test cases. Based on an initial human analysis, we observe that in several cases, the model begins with a correct sequence but introduces a critical error—often in the form of incorrect code structures—at a specific step. For example, instead of generating a for loop to iterate over remaining elements, the model might incorrectly rank the return token at the top position, prematurely terminating the control flow. At these error points, the correct next token is not missing from the candidate set but is ranked below the top-1 position (commonly second or third), resulting in an irreversible deviation from the correct logic. This suggests that some generation failures are caused by local ranking errors at specific decoding steps, ultimately leading to incorrect code behavior. We refer to such steps as logic drift points during decoding and aim to identify them.
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. For instance, a structural discrepancy is flagged when:
  1. (1)
    control-flow node types differ (e.g., if vs. for);
  2. (2)
    relational or boolean operators differ (e.g., > vs. >=);
  3. (3)
    loop or branching boundaries diverge;
  4. (4)
    indentation-aligned block structures differ.
Any identified mismatch is traced back to its corresponding line in the original code and extracted as a candidate drift point.
For each candidate line, we further determine whether the drift occurs at the beginning of the line and record this for statistical analysis. Finally, each candidate is reviewed by a human annotator to confirm the validity of the drift and to annotate the precise start token. To ensure consistency, annotators adhered to the following explicit guidelines:
  1. (i)
    focus strictly on structural (AST-level) differences;
  2. (ii)
    ignore lexical or superficial variations (e.g., variable names, comments);
  3. (iii)
    label the earliest divergence point;
  4. (iv)
    treat alternative valid implementations as non-drift unless the semantics fundamentally change.
For instance, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
@@ 2.3.1. Method
We apply this next-token prediction procedure to a range of open-source LLMs, including five code-specific models, DS-1.3B, DS-6.7B, ST-3B, QW2.5-1.5B, and QW2.5-7B, as well as four general-purpose models: QW3-0.6B, QW3-1.7B, QW3-4B, and QW3-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, and 129,165 for Qwen series.
@@ 2.3.2. Results
Model DS-1.3B DS-6.7B ST-3B QW2.5-1.5B QW2.5-7B QW3-0.6B QW3-1.7B QW3-4B QW3-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.4906 0.4755 0.5249 0.5028 0.4936 0.4877
Figure 4 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 4 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1.2, QW3-0.6B has approximately 13% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW3-1.7B) exceed a value of 5.5, and the rest fall in the range of 3.5 to 4.7.
@@ 3.2.1. Data Collection
We cast threshold learning as a binary classification task. Each decoding step is a sample, where the input feature is the entropy $H(\bm{p}_{t})$ and the label is 1 if the ground-truth token is the top-1 prediction, and 0 otherwise. Training data is generated via ground-truth–guided next-token prediction on BigCodeBench (Zhuo et al., 2025). To strictly mitigate data contamination, we verified semantic similarity between BigCodeBench and our evaluation sets using SBERT (Reimers and Gurevych, 2019). Manual review of all problem pairs exceeding a 0.7 cosine similarity threshold (maximum observed was 0.75) confirmed zero overlapping duplicates. Through this process, we collect approximately 100K––160K positive samples and 16K–23K negative samples per model (e.g., 161K positive and 19K negative for DS-6.7B).